OpenStudy (anonymous):

Solve the equation X=7/8=4

OpenStudy (anonymous):

do u mean x=7/x=4

OpenStudy (anonymous):

@alanaleigh13

OpenStudy (anonymous):

X+7/8=4

OpenStudy (anonymous):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@alanaleigh13 do you understand?

OpenStudy (anonymous):

nope and thats not an answer

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

their like this a: x= 4 and 7/8 like a mixed fraction